# A Roadmap for Simulating Chemical Dynamics on a Parametrically Driven Bosonic Quantum Device

### v1

@@ 1.2 Computational Methods
The matrix exponential operator is implemented using the scipy.linalg.expm routine, which implements a scaling and squaring algorithm based on Pade’s approximation. 49
@@ 1.3 Initial State
where $x$ indicates the position, $x_{0}$ indicates the cutoff position and $t$ the smoothness of the function near the cutoff. We observe that the Heaviside function, $\Theta(x;x_{0})=1$ if $x\geq x_{0}$ else $\Theta(x;x_{0})=0$ , is recovered when taking $\lim_{t\rightarrow 0}S(x;x_{0},t)$. This allows localization of the initial state in position space which we then convert it back to the Fock basis representation (figure 7, right panel). All dynamics trajectories use eigenstate selection with a sigmoidal filter with a tail of 0.5.
@@ 1.4 Dynamics Subspace
The accuracy of the dynamics is dependent on the number of Fock basis states used. However, the size of the Lindbladian matrix scales as $\mathcal{O}(N^{4})$ with the number of Fock states as compared to the Hamiltonian $\mathcal{O}(N^{2})$, which makes matrix exponentiation (performed once for each set of Hamiltonian parameters) and multiplication (performed for each timestep in a trajectory) a limiting factor in simulation. Thus, we generate the complete Hamiltonian with a large number of Fock states (N=300) and numerically diagonalize to obtain accurate eigenvalues and eigenvectors and use the first M=20 states to perform as similarity transformation matrix to reduce the dynamics computational space:
@@ 1.5 Observables
2 Simulating barrier crossing dynamics on a Kerr-Cat device
@@
    As the inverse of the real value of the smallest singular value of the Lindbladian supermatrix, obtained via diagonalization, as in figure 8a.
@@
The two methods for determining $T_{X}$ gave similar results and found to exhibit the same behavior when it comes to the dependence of $T_{X}$ on the Kerr-Cat parameters. The results reported in the text were obtained via method 1 unless otherwise noted (figure 8b).
@@
Figure 8: Relaxation timescales associated with the (a) Lindbladian maximal real eigenvalue (left) and (b) the dynamical relaxation rates (right) obtained by fitting the population traces as a function of the Kerr-Cat parameters $\varepsilon_{1}$ and $\varepsilon_{2}$. Both plots use dissipation parameters $\kappa=0.1$, $n_{\text{th}}=0.1$.
The dependence of $T_{X}$ on the Kerr-Cat parameters $\varepsilon_{1}$ and $\varepsilon_{2}$ is shown in Fig. 8. The plot shows a rich structure including (a) a zone in the lower right corner where the barrier crossing is very fast, which corresponds to a low barrier or a complete lack of a barrier, (b) Fast barrier crossing in the upper left corner for particular values of ($\varepsilon_{1},\varepsilon_{2}$) where the energy levels in the reactant and product wells are in resonance (see white lines in Figs. 8 and SI), and (c) alternation between ”broad” and ”thin” resonance transitions both as a function of $\varepsilon_{1}$ for fixed $\varepsilon_{2}$ and along the ($\varepsilon_{1},\varepsilon_{2}$) resonance line.
@@
The fast barrier crossing regimes correspond to resonances between the energy levels in the reactant and product wells which lead to efficient tunneling through the barrier. The aforementioned ”thin” regions correspond to eigenstate overlap close the top of the barrier as well as a high state density at the barrier top, providing a transient state to retain population before decaying to the global ground state (see figure 8 and corresponding panels S1, 2, 3 of figure 9). By contrast the ”broad” regimes contain degenerate states but no density centered at the top of the barrier, thus reducing the overall population transfer rate (see figure 8 and corresponding panels S2, 1, 4 of figure 9). However, the rate is nonetheless enhanced due to the presence of quasi-degenerate states near the barrier top, providing a pathway for barrier crossing. Furthermore, the intermediate states without resonant states have longer lifetimes due to lack of overlap between eigenstates or localized state density on either side of the well (see figure 8 and corresponding panels N1 and N2 of figure 9). Going along each of the resonance lines showcases an increasing number of degenerate pairs of states with increasing $\varepsilon_{2}$ (figure 8: S1-S2; 1-2; 3-4). Going between different resonance lines with increasing $\varepsilon_{1}$ changes the first state index in resonance between the two wells (S1: first and second, 1: second and third, 3: third and fourth).
@@
Finally, Fig. 11 shows how the dependence of $T_{X}$ on $\epsilon_{1}$ and $\epsilon_{2}$ is impacted by the bath temperature, as measured by $n_{th}$. While $T_{X}$ shows an overall increase with decreasing $n_{th}$, the structure is seen to be minimally impacted by changing $n_{th}$.
3 Basis Set Convergence
In this section, we explore the convergence of the Kerr-Cat Hamiltonian parameters to the number of Eigen-basis used. In Fig. 12, we present the timescales obtained from the exact diagonalization of the Lindbladian as a function of Hamiltonian parameters $\varepsilon_{1}$ and $\varepsilon_{2}$, for a different number of Eigen-basis. As can be appreciated, with $n_{basis}>10$ the timescales are semi-quantitative converged, and with $n_{basis}>20$ quantitative agreement is found. Unless otherwise stated, we used $n_{basis}=20$ for the device dynamics.
@@ 4 Chemical Potentials
Thymine-Adenine (DNA)37 0.02068986 0.00525515 0.0413797 0.0157655 $E_{h}$
Malonaldehyde (cis-trans)39, 40 0.00009374 0.000109 0.00299 0.005232 a.u.
Malonaldehyde (cis-cis)39, 40 0.000714286 0 0.004 0 a.u.
@@
Furthermore, we note that the adenine-thymine potential is expressed in terms of a unitless length parameter $\zeta=x/x_{0}$, which has been estimated to be $x_{0}=1.9592\,a_{0}$ based on matching the energy eigenvalues listed in reference 37. For the dynamics shown in this work, the given literature potentials where fit to the simpler double-well potential, without a cubic polynomial term,
@@
5 Dynamics Trajectories and Rate Fits
@@
6 Basis Benchmark for Dynamical Evolution
This section showcases the dynamics evolution figures for the four systems, using a different number of basis set to showcase the convergence of the results listed in the main text. The figures showcase the results of using a Fock space of dimension 50, 100 and 150 to accurately represent the Hamiltonian and eigenstates, which are then used for the dynamics propagation in a reduced subspace. For all cases, 50 eigenstates were used for the dynamics propagation, which incorporate many levels beyond the barrier top energy. For some of the listed systems, dynamics convergence with respect to number of eigenstates is observed for values much smaller than than fifty. Deviations of the Kerr-Cat fitted rates at larger c-values, result from the Hamiltonian perturbative terms that are sensitive to the value of $c$ and errors associated with the fitting protocol.

### v2

@@ 1.2 Computational Methods
The matrix exponential operator is implemented using the scipy.linalg.expm routine, which implements a scaling and squaring algorithm based on Pade’s approximation. 48
@@ 1.3 Initial State
where $x$ indicates the position, $x_{0}$ indicates the cutoff position and $t$ the smoothness of the function near the cutoff. We observe that the Heaviside function, $\Theta(x;x_{0})=1$ if $x\geq x_{0}$ else $\Theta(x;x_{0})=0$ , is recovered when taking $\lim_{t\rightarrow 0}S(x;x_{0},t)$. This allows localization of the initial state in position space which we then convert back to the Fock basis representation (figure 7, right panel). All dynamics trajectories use eigenstate selection with a sigmoidal filter with a tail of 0.5.
@@ 1.4 Dynamics Subspace
The accuracy of the dynamics is dependent on the number of Fock basis states used. However, the size of the Lindbladian matrix scales as $\mathcal{O}(N^{4})$ with the number of Fock states as compared to the Hamiltonian $\mathcal{O}(N^{2})$, which makes matrix exponentiation (performed once for each set of Hamiltonian parameters) and multiplication (performed for each timestep in a trajectory) a limiting factor in simulation. Thus, we generate the complete Hamiltonian with a large number of Fock states (N=300) and numerically diagonalize to obtain accurate eigenvalues and eigenvectors and use the first M=20 states to perform a similarity transformation matrix to reduce the dynamics computational space:
@@ 1.5 Observables
1.6 Simulating barrier crossing dynamics on a Kerr-cat device
@@
    As the inverse of the maximum amplitude non-zero real eigenvalue of the Lindbladian supermatrix, obtained via diagonalization, as in figure 8a.
@@
The two methods for determining $T_{X}$ gave similar results and were found to exhibit the same behavior when it comes to the dependence of $T_{X}$ on the Kerr-cat parameters. The results reported in the text were obtained via method 1 unless otherwise noted (figure 8b).
@@
Figure 8: Relaxation timescales associated with the (a) Lindbladian maximal real eigenvalue (left) and (b) the dynamical relaxation rates (right) obtained by fitting the population traces as a function of the Kerr-cat parameters $\epsilon_{1}$ and $\epsilon_{2}$. Both plots use dissipation parameters $\kappa=0.1$, $n_{\text{th}}=0.1$.
The dependence of $T_{X}$ on the Kerr-cat parameters $\epsilon_{1}$ and $\epsilon_{2}$ is shown in Fig. 8. The plot shows a rich structure including (a) a zone in the lower right corner where the barrier crossing is very fast, which corresponds to a low barrier or a complete lack of a barrier, (b) Fast barrier crossing in the upper left corner for particular values of ($\epsilon_{1},\epsilon_{2}$) where the energy levels in the reactant and product wells are in resonance (see white lines in Figs. 8 and SI), and (c) alternation between “broad” and “thin” resonance transitions both as a function of $\epsilon_{1}$ for fixed $\epsilon_{2}$ and along the ($\epsilon_{1},\epsilon_{2}$) resonance line.
@@
The fast barrier crossing regimes correspond to resonances between the energy levels in the reactant and product wells which lead to efficient tunneling through the barrier. The aforementioned “thin” regions correspond to eigenstate overlap near the top of the barrier as well as a high state density at the barrier top, providing a transient state to retain population before decaying to the global ground state (see figure 8 and corresponding panels S1, 2, 3 of figure 9). By contrast the “broad” regimes contain degenerate states but no density centered at the top of the barrier, thus reducing the overall population transfer rate (see figure 8 and corresponding panels S2, 1, 4 of figure 9). However, the rate is nonetheless enhanced due to the presence of quasi-degenerate states near the barrier top, providing a pathway for barrier crossing. Furthermore, the intermediate states without resonant states have longer lifetimes due to lack of overlap between eigenstates or localized state density on either side of the well (see figure 8 and corresponding panels N1 and N2 of figure 9). Going along each of the resonance lines showcases an increasing number of degenerate pairs of states with increasing $\epsilon_{2}$ (figure 8: S1-S2; 1-2; 3-4). Going between different resonance lines with increasing $\epsilon_{1}$ changes the first state index in resonance between the two wells (S1: first and second, 1: second and third, 3: third and fourth).
@@
Finally, Fig. 11 shows how the dependence of $T_{X}$ on $\epsilon_{1}$ and $\epsilon_{2}$ is impacted by the bath temperature, as measured by $n_{\text{th}}$. While $T_{X}$ shows an overall increase with decreasing $n_{\text{th}}$, the structure is seen to be minimally impacted by changing $n_{\text{th}}$.
1.7 Basis Set Convergence
In this section, we explore the convergence of the Kerr-cat Hamiltonian parameters to the number of Eigen-basis used. In Fig. 12, we present the timescales obtained from the exact diagonalization of the Lindbladian as a function of Hamiltonian parameters $\epsilon_{1}$ and $\epsilon_{2}$, for a different number of Eigen-basis. As can be appreciated, with $n_{basis}>10$ the timescales are semi-quantitative converged, and with $n_{basis}>20$ quantitative agreement is found. Unless otherwise stated, we used $n_{basis}=20$ for the device dynamics.
@@ 1.8 Chemical Potentials
Thymine-Adenine (DNA)38 0.02068986 0.00525515 0.0413797 0.0157655 $E_{h}$
Malonaldehyde (cis-trans)40, 41 0.00009374 0.000109 0.00299 0.005232 a.u.
Malonaldehyde (cis-cis)40, 41 0.000714286 0 0.004 0 a.u.
@@
Furthermore, we note that the adenine-thymine potential is expressed in terms of a unitless length parameter $\zeta=x/x_{0}$, which has been estimated to be $x_{0}=1.9592\,a_{0}$ based on matching the energy eigenvalues listed in reference 38. For the dynamics shown in this work, the given literature potentials where fit to the simpler double-well potential, without a cubic polynomial term,
@@
1.9 Dynamics Trajectories and Rate Fits
@@
1.10 Basis Benchmark for Dynamical Evolution
This section showcases the dynamics evolution figures for the four systems, using a different number of basis set to showcase the convergence of the results listed in the main text. The figures showcase the results of using a Fock space of dimension 50, 100 and 150 to accurately represent the Hamiltonian and eigenstates, which are then used for the dynamics propagation in a reduced subspace. For all cases, 50 eigenstates were used for the dynamics propagation, which incorporate many levels beyond the barrier top energy. For some of the listed systems, dynamics convergence with respect to number of eigenstates is observed for values much smaller than fifty. Deviations of the Kerr-cat fitted rates at larger $c$-values result from the Hamiltonian perturbative terms that are sensitive to the value of $c$ and errors associated with the fitting protocol.